OpenStudy (anonymous):

Your friend hands you a graph of the performance statistics of the newest car to roll off the assembly line. He says, “I know this graph is f(x) = –4(x – 5)3 + 7 but I can’t remember how it is related to the graph of x3.” Explain to your friend how the graph f(x) is a translation of the graph x3.

OpenStudy (anonymous):

someone help me?

OpenStudy (anonymous):

what do you mean by the 1st sentence? qhere did you get x^4?

OpenStudy (jchick):

Sorry wrong equation

OpenStudy (anonymous):

im confused now...

OpenStudy (jchick):

Sorry i was trying to post in a different one

OpenStudy (jchick):

I will attempt to help you though

OpenStudy (jchick):

Hang on and let me work this out

OpenStudy (anonymous):

ok

OpenStudy (jchick):

y=(x)^3 Let's try to turn this into the performance function. Putting a negative in front flips it over the x-axis. y=−(x)^3 To make shifts to the right or up we subtract that value from the corresponding variable. So we want to shift this function to the right 5 units, so we'll subtract 5 from x. y=−(x−5)^3 We also want to shift this function up by 7, so we'll subtract 7 from y. y−7=−(x−5)^3 Then add 7 to each side, y=−(x−5)^3+7 Then we'll squeeze the function (make it grow four times faster) by multiplying it by 4. y=−4(x−5)^3+7

OpenStudy (anonymous):

oooooh i get it!! thank you so much!

OpenStudy (jchick):

No problem

OpenStudy (anonymous):

can you help me with another question?

OpenStudy (jchick):

Sure what is it?

OpenStudy (anonymous):

In your lab, a substance's temperature has been observed to follow the function T(x) = (x - 4)3 + 6. The turning point of the graph is where the substance changes from a liquid to a gas. Explain to your fellow scientists how to find the turning point of this function, using complete sentences.

OpenStudy (jchick):

I am actually having some trouble doing this one

OpenStudy (jchick):

Sorry

OpenStudy (jchick):

Any others?

OpenStudy (anonymous):

its ok, thaks for trying though

OpenStudy (jchick):

yup no problem

OpenStudy (jchick):

Do you have any more questions?

OpenStudy (anonymous):

no, thanks

OpenStudy (jchick):

Ok are you good with graphs?